OpenStudy (anonymous):

a line is drawn through point (1,2) forming a right triangle with the positive x- and y-axes. The slope of the line forming the triangle of least area is...

OpenStudy (anonymous):

got anything??

OpenStudy (anonymous):

not yet lol

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

but im making dinner for my kids and thinking it over

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

are you expected to use calculus or vectors to solve this one?

OpenStudy (anonymous):

calc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its a question from my practice AP calc book

OpenStudy (anonymous):

i know the answer, i just don't know how to get it

OpenStudy (anonymous):

yeah, that's always the hard part hehe

OpenStudy (anonymous):

always! haha

OpenStudy (anonymous):

what part of calc are you in? multivariable, langrange multipliers or just single variable?

OpenStudy (anonymous):

uhmm i'm in Calc AB..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok. Well the area of this right triangle is:\[A=\frac{1}{2}xy\]Now we need to relate x and y with one another. The line through the point (1,2) is our hypotenuse. It crosses the x axis at the point (0,y) and it crosses the y-axis at the point (x,0). The slope of this line is:\[m=\frac{y-2}{x-1}\]Taking the point (x,y)=(0,y) we get \[y=2-m\] where m is a constant. We have:\[A(x)=\frac{1}{2}(x)(2-m)\]Now we differentiate and set equal to zero to find the critical points:\[A'(x)=\frac{2-m}{2}=0\]so, m=2.

OpenStudy (anonymous):

hmm. well my answer key says its supposed to be -2..

OpenStudy (anonymous):

im missing something here. It makes sense that the slope is negative though since the triangle is bound by the x and y axes.

OpenStudy (anonymous):

yeah, that's why i can't figure it out either lol

OpenStudy (anonymous):

I know its a simple matter of a negative sign but I cant justify it!?!

OpenStudy (anonymous):

sorry, if I figure it out later Ill let you know but too many distractions tonight!

OpenStudy (anonymous):

alright, well thanks for trying! :)